OpenStudy (legomyego180):

Greens Theorum:

OpenStudy (legomyego180):

Integrate h(x, y) = yi - xj over the triangle with vertices (-1, 0), (1, 0) and (0, 1) traversed counterclockwise.

OpenStudy (legomyego180):

\[\oint (\frac{ dP }{ dy }-\frac{ dQ }{ dx })dA\] or something like this. So I set up: \[\int\limits\limits_{-1}^{0}\int\limits\limits_{0}^{x}-1-1dydx\]

OpenStudy (legomyego180):

The answer is -1 but im coming up with -1

zepdrix (zepdrix):

lol what?

OpenStudy (legomyego180):

lol greens theorum zep. comon! I dropped a negative, should be 1. Anyway its asking me to integrate this triangle... just figured it out I had my bounds wrong. Thanks zep!

zepdrix (zepdrix):

oh ok XD

OpenStudy (legomyego180):

wait actually, what would the y bounds be for my double integral if i was integrating this triangle? 

zepdrix (zepdrix):

Hmm I was thinking of it like this I guess... In the x-direction you're bounded by the curves x=-y on the left, and x=y on the right. \(\rm -y\le x\le y\) And then your y is bound by the bottom and top, \(\rm 0\le y\le1\) But ummm.. you can probably do it your way... lemme see if that makes sense..

zepdrix (zepdrix):

Oh yes, if you try to do it in the other direction, you run into a problem right? Because you have to split it into two separate integrals. The upper boundary `changes` in the middle! :O

OpenStudy (legomyego180):

With the symmetry I could just multiply it by two and integrate from 0 to 1 or -1 to 0 I think then multiply by 2, but that still would get me a different answer.

zepdrix (zepdrix):